Chuck and Dana agree to meet in Chicago for the weekend. Chuck travels 147 miles in the same time that Dana travels 129 miles. If Chuck's rate of travel is 6 mph more that Dana's, and they travel the same length of time, at what speed does chuck travel?
Answer by mananth(16946)

Dana x 129 miles
Chuck x+ 8 147 miles
The time taken by them is the same
t=d/r
129 / x = 147 /(x+ 8 )
Cross multiply
129 (x + 8 ) = 147 * x
129 x + 1032 = 147 x
147 x -129 x = 1032
18 x = 1032
/ 18
x = 57.33 Dana mph
65.33 Chuck mph